The household supply of electricity is at 220 V (rms value) and 50 Hz. Find the peak voltage and the least possible time in which the voltage can change crom the rms value to zero.

Grade:11

2 Answers

Navjyot Kalra
9 years ago
Sol. E baserms = 220 V Frequency = 50 Hz (a) E base rms = E base 0/√2 ⇒ E base 0 = E base rms √2 =√2 * 220 = 1.414 * 220 = 311.08 V = 311 V (b) Tiem taken for the current to reach the peak value = Time taken to reach the 0 value from r.m.s I = I base 0/√2 ⇒ I base 0/√2 = I base 0 Sin ωt ⇒ ωt = π/4 ⇒ t = π/4∗ = π/4 *2πf = π/8π50 = 1/400 = 2.5 ms
Arun Kumar IIT Delhi
9 years ago
Hello
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty